OpenStudy (anonymous):

The length of a rectangle is 3mm more than twice its width, and the area of the rectangle is 77mm^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

length = l width = w l=2w+3 area = l*w \[lw=2w^{2}+3w\] but area is 77 \[77=2w^{2}+3w\]

OpenStudy (pradipgr817):

length x (2 x length +3) = 77 sole it

OpenStudy (anonymous):

sole the quadratic equation find possible value of w and substitute that value in other equation to get l

OpenStudy (anonymous):

solve*

OpenStudy (anonymous):

im lost sorry, this is not my thing

OpenStudy (anonymous):

let length be L,width be W  it is give that length is 3mm more than twice the width..... think like this... twice the width is .......2W 3mm more than twice the width is .......3+2W 3mm more than twice the width IS LENGTH ......L so L=3+2W.. but AREA= length * width=L*W but we got L = 3+2W so AREA = (3+2W)*W...expand it area is give as 77 77= 3W+2W^2 \[77=3W+2W^{2}\]